In Pleasantville about 25% of the residents prefer vanilla ice cream over other flavors. A certain day, 175 residents of Pleasantville went to an ice cream store, what is the probabil- ity that 50 or more will get vanilla ice cream?


1
Expert's answer
2021-05-07T14:17:34-0400

p = 0.25

q = 0.75

n = 175

Using binomial distribution:

Mean:

"\mu = n \times p = 175 \times 0.25 = 43.75"

Standard deviation:

"\sigma = \sqrt{n \times p \times q} = \sqrt{175 \times 0.25 \times 0.75} = 5.728"

"P(X≥50) = 1 -P(X<49.5) \\

= 1 -P(Z< \frac{49.5 -43.75}{5.728}) \\

= 1 -P(Z<0.197) \\

= 1 -0.5780 \\

= 0.422"